# RDMA-Based Algorithms for Sparse Matrix Multiplication on GPUs

### v1

@@ Abstract
I Introduction
@@
Sparse matrix multiplication is an important computational primitive that arises in simulation, data analysis, and machine learning applications. Typically limited by memory and network performance, these computations are especially challenging for unstructured matrices, such as those occurring in graph neural networks, genomics, graph analytics, and other data intensive problems. Sparse matrix primitives provide a convenient and important set of operations for performance tuning that will impact a wide array of applications, and there is a large body of prior work optimizing sparse matrix kernels for multicore [1], manycore [2], GPU [3], and distributed memory environments [4, 5, 6, 7, 8]. Two of the most important kernels are sparse times dense matrix multiplication (SpMM), where the dense matrix is tall and skinny, i.e., representing a set of vectors, and sparse times sparse matrix multiplication (SpGEMM). Sparse matrix-vector multiplication is also an important kernel in many applications, and has been studied even more extensively, although our focus here is on the matrix-matrix operations.
@@
The large majority of prior work on distributed-memory sparse matrix kernels focuses on bulk synchronous implementations based on SUMMA [9], using collective operations, typically broadcast and/or reduce, in the inner loop of the algorithm to move data to the processes that need it for their local matrix multiplications. While SUMMA-like algorithms have the advantages of being straightforward to implement and generally scaling well, for sparse problems, they have the disadvantage of forcing processes to proceed in lockstep in the inner loop of the algorithm. When there is load imbalance, either in the amount of computation that must be performed locally by each process, the amount of data that must be transferred to each process, or both, this can potentially result in decreased performance. This can occur even when the total amount of work performed by each process is the same, since processes may have differing amounts of work in each iteration. To illustrate, Figure 1 shows two kinds of load balance for squaring a randomized sparse matrix generated by the R-MAT model [10], with parameters $a=0.6$, $b=c=d=0.4/3$, edgefactor $8$, and scale $17$. The total (end-to-end) computation has only $20\%$ load imbalance whereas the synchronization points in between stages amplify the load imbalance to 2.3$\approx$ 2.3$\times$ 2.3 ×. The load imbalance is defined as the max/avg ratio, the ratio of maximum number of flops performed by any processor to the average number of flops per processor.
@@
Load imbalance is often addressed by randomly permuting the rows and/or columns of the sparse matrices, but this has disadvantages, the foremost being that the required permutation of the inputs, followed by permutation of the outputs, can be expensive. Furthermore, permuting the matrix may possibly remove structural locality, resulting in decreased performance of local matrix multiply operations. For example, Slota et al. [11] showed that the loss of locality due to random permutations can hurt the performance of graph algorithms significantly where parallel PageRank performs 25×2-5$\times$2 - 5 × slower on a randomly permuted graph compared to other orderings. Even if we ignore the loss of locality or try to recover it by local reordering after a global random permutation, there are theoretical limits on how much load balance random permutations can achieve. For matrices with dense rows or columns, the load balance can deteriorate with $\sqrt{p}$ for large $p$, where $p$ is the number of processors in the worst case, per Theorem 5.4 of Azad et al. [12].
@@
Figure 1: Total (end-to-end) vs.  per-stage load balance multiplying a R-MAT model-generated sparse matrix with a sparse 2D algorithm. Simulated on a $16\times 16$ process grid.
@@
II Background
Sparse times dense matrix multiply (SpMM) and sparse times sparse matrix multiply (SpGEMM) are two important sparse linear algebra primitives. SpMM is used in a variety of blocked iterative methods, graph algorithms, as well as graph neural networks (GNNs) [13, 14, 15]. In these contexts, SpMM typically involves multiplication of a sparse matrix with a tall and skinny dense matrix, where the number of columns of the sparse matrix varies between 32 and 1024. SpGEMM is also used in graph algorithms, including betweenness centrality, triangle counting, cycle detection, and Markov clustering, as well as in applications such as in genomics [16, 17, 18, 19, 20].
II-A Distributed Matrix Algorithms
Distributed matrix algorithms allow multiple processes to work together to execute a single matrix multiply operation. Let’s first consider a distributed matrix multiply computing the product $C=AB$. Here, each of the matrices $C$, $A$, and $B$ is split up into tiles which live on different processes. The matrix tiles will be sparse or dense, depending on what kind of matrix multiplication is being performed. Let’s assume that $C$ is distributed among a grid of $M\times N$ tiles, $A$ is distributed among a grid of $M\times K$ tiles, and $B$ is distributed among a grid of $K\times N$ tiles. The algorithms we consider in this paper rely on a tiling scheme that is regular on the indices. In other words, each tile has the same dimensions, modulo differences due to the tile size not dividing the matrix size evenly. A different, more irregular tiling mechanism might be more appropriate for matrices from certain applications such as those involving linear-scaling quantum-chemical calculations [21], where the near-sightedness creates natural block-sparse structure. However, the vast majority of application targets (GNNs, graph algorithms, eigenvector computations, and tensor decompositions) do not have this natural block-sparse structure, so the added complexity of irregular tiling is not justified for us.
@@ II-A Distributed Matrix Algorithms
II-B Bulk Synchronous SUMMA
@@
As discussed in Section I, variation in the number of nonzeros, as well as the sparsity patterns, of individual tiles of sparse input matrices can create load imbalance in computation due to the differing numbers of flops that must be performed in each local matrix multiply, as well as load imbalance in communication, due to the different amounts of data that must be transferred by each process. Bulk synchronous SUMMA-like algorithms can suffer due to these load imbalances.
II-C RDMA and Asynchrony
Remote Direct Memory Access (RDMA) is a hardware capability offered by the Network Interface Cards (NICs) in most modern supercomputers and datacenters. Using RDMA, a process can expose a region of its memory, often referred to as “shared memory,” to be directly accessible by remote processes. Processes can then issue put, get, and atomic operations to remotely access and manipulate the shared memory regions of other processes. These RDMA operations are handled directly by the NIC on the remote node, hence no coordination is required in order to access the memory of a remote process. Some of the commonly used frameworks that support direct RDMA access include OpenSHMEM, GASNet-EX, and MPI’s one-sided communication API. For distributed programs that run on GPUs, systems with Infiniband can currently take advantage of GPUDirect RDMA, which allows direct transfers from GPU-to-GPU over the network. NVSHMEM, an extension of OpenSHMEM from NVIDIA, allows for direct transfers of data between GPUs, either within a node over PCIe or NVLink, or between nodes using GPUDirect RDMA over Infiniband. In this paper, we use NVSHMEM to directly copy between GPUs. A number of different libraries have been developed that use RDMA primitives to build data structures and algorithms, with some focusing particularly on irregular data structures and algorithms, where asynchronous RDMA primitives can be particularly beneficial, both for performance and software engineering reasons [22, 23, 24].
III RDMA-Based Algorithms
III-A Data Structures
@@ III-A1 Reading Tiles
III-A2 Modifying Remote Tiles
@@
III-B Algorithms
III-B1 RDMA Stationary C Algorithm
The most straightforward RDMA-based algorithm is the stationary C algorithm, in which each processor iterates through its tiles of the C matrix, and, for each tile, iterates through the corresponding row block of A and column block of B, retrieving each of the remote tiles via get_tile and multiplying them together using a local matrix multiply operation. Since the method get_tile does not require any coordination with remote processors, each process can execute its work independently, and does not need to wait on or synchronize with other processors. In order to achieve good performance, a few optimizations are required, including prefetching tiles for the next iteration, which allows overlap of computation and communication, and an iteration offset in the inner loop, which balances which tile is requested within each row and column and generally ensures the first RDMA get is to the local tile. These optimizations are discussed in detail in Section III-C.
for i in 0..M-1:
for k in 0..K-1:
%**** sparse-gpu.tex Line 350 ****if A.owner(i, k) == rank():
for j in 0..N-1:
local_a = A.tile_ref(i, k)
local_b = B.get_tile(k, j)
local_c = local_a*local_b
queue[C.owner(i, j)].push(get_ptr(local_c))
[…]
C.my_tile() += get_tile(local_c_ptr)
barrier()
@@ III-B1 RDMA Stationary C Algorithm
III-B2 RDMA Stationary A and B Algorithms
@@ III-C Optimizations
for i in 0..M-1:
for j in 0..N-1:
if C.owner(i, j) == rank():
k_offset = i + j
buf_a = A.async_get_tile(i, k_offset %K)buf_b = B.async_get_tile(k_offset %K, j)for k_ in 0..K-1:
k = (k_ + k_offset) %Klocal_a = buf_a.get()
local_b = buf_b.get()
local_c = C.tile_ref(i, j)
[…]
buf_a = A.async_get_tile(i, (k+1) %K)buf_b = B.async_get_tile((k+1) %K, j)local_c += local_a*local_b
barrier()
@@ III-D Workstealing Algorithms
In locality-aware workstealing, instead of stealing randomly, each process will only steal pieces of work where it owns one of the components, ensuring that the cost of performing stolen work will be similar to if it had been performed by the origin process. Locality-aware workstealing requires a 3D work grid, where element $i,j,k$ corresponds to the component matrix multiply $C[i,j]+=A[i,k]*B[k,j]$. Processes will first perform work associated with the original stationary distribution, and then attempt to steal work items for which they have one of the three components.
def attempt_work(i, k):
# Remote atomic fetch-and-add to reserve work
my_j = reserve_grid[i, k]++
while my_j < N:
local_a = A.get_tile(i, k)
local_b = B.get_tile(k, j)
local_c = local_a*local_b
[…]
my_j = reserve_grid[i, k]++
[…]
for i in 0..M-1:
for k in 0..K-1:
if A.owner(i, k) == rank():
attempt_work(i, k)
[…]
for idx in 0..M-1:
i = (rank()+idx) / M
k = (rank()+idx) %Mattempt_work(i, k)
[…]
C.my_tile() += get_tile(local_c_ptr)
barrier()
@@ IV Performance Model
If we offset the order of iteration of the $k$ loop by $i+j$, as discussed in Section III-C, we can also ensure that communication will be balanced with each process sending exactly two tiles per iteration, one of the A matrix, and one of the B matrix. To prove this, assume tile $(i,j)$ is stored on processor $i(N+1)+j$, where $i\in[0,N)$, $j\in[0,N)$, and $N=\sqrt{p}$.
@@
Figure 2: Inter-node roofline plots for SpMM and SpGEMM with a 2D distribution. SpMM plot models performance for different widths of the dense B matrix at a fixed scale (24 GPUs), while SpGEMM models performance at different scales. Dashed horizontal lines represent local roofline peaks for SpMM and SpGEMM operations, while vertical lines represent inter-node roofline peaks for particular problems.
@@
Note that the roofline model is dependent on the sparsity structure, since $cf$ depends on the particular sparse matrices being multiplied. Our inter-node SpGEMM roofline model is plotted in Figure 2. To obtain realistic values for cf and the number of flops performed FLOPS(A, B), we performed each of the component local SpGEMM operations in the distributed SpGEMM operation for the isolates subgraph4 matrix, recording their values for cf and FLOPS(A, B) to compute average values for different numbers of GPUs $P$. In the SPGEMM plot, inter-node roofline peaks are much closer to their local roofline peaks than in the SpMM plot. Thus, our roofline model suggests SpGEMM is significantly less network communication-bound than SpMM, although still communication bound. We discuss more conclusions from our roofline model in Section VI.
V Implementation
V-A Communication Layer
As discussed in Section II-C, the current generation of GPU-based supercomputers offers GPUDirect RDMA over Infiniband networks, which ensures that issuing one-sided operations will be efficiently executed in hardware. In GPUDirect RDMA, the CPU prepares an Infiniband request, which it sends to the local network interface card (NIC). When the NIC receives the request, it will copy the data directly over the network from GPU to GPU, without staging data through the CPU. When copies take place between GPUs within the same node, the high-bandwidth NVLink fabric is used.
@@ V-B Data Structures
We implement sparse and dense distributed matrix data structures using BCL, an RDMA-based distributed data structures library in written in C++. Both dense and sparse matrix data structures split the distributed matrix up into evenly sized tiles, which are then assigned to processors using a processor grid, as in ScaLAPACK. As discussed in Section III-A, each process has a copy of a directory of remote pointers, which point to the matrix tiles stored in pinned memory, directly accessible through RDMA operations using NVSHMEM. In the dense case, a single remote pointer points to a dense tile data structure, and in the sparse case three remote pointers point to the values, row pointer, and column index arrays of a compressed sparse row (CSR) data structure. All data is stored directly on GPUs, and can be copied directly from GPU to GPU over the network using BCL’s NVSHMEM backend.
V-C Data Access Primitives

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 38th ACM International Conference on Supercomputing; June 4–7, 2024; Kyoto, Japanbooktitle: Proceedings of the 38th ACM International Conference on Supercomputing (ICS ’24), June 4–7, 2024, Kyoto, Japandoi: 10.1145/3650200.3656623isbn: 979-8-4007-0610-3/24/06
1. Introduction
@@
Sparse matrix multiplication is an important computational primitive that arises in simulation, data analysis, and machine learning applications. Typically limited by memory and network performance, these computations are especially challenging for unstructured matrices, such as those occurring in graph neural networks, genomics, graph analytics, and other data intensive problems. Sparse matrix primitives provide a convenient and important set of operations for performance tuning that will impact a wide array of applications, and there is a large body of prior work optimizing sparse matrix kernels for multicore (Patwary et al., 2015), manycore (Saule et al., 2013), GPU (Yang et al., 2018), and distributed memory environments (Schubert et al., 2011; Acer et al., 2016; Hong et al., 2019; Solomonik and Hoefler, 2015; Gu et al., 2020). Two of the most important kernels are sparse times dense matrix multiplication (SpMM), where the dense matrix is tall and skinny, i.e., representing a set of vectors, and sparse times sparse matrix multiplication (SpGEMM). Sparse matrix-vector multiplication is also an important kernel in many applications, and has been studied even more extensively, although our focus here is on the matrix-matrix operations.
@@
The large majority of prior work on distributed-memory sparse matrix kernels focuses on bulk synchronous implementations based on SUMMA (Van De Geijn and Watts, 1997), using collective operations, typically broadcast and/or reduce, in the inner loop of the algorithm to move data to the processes that need it for their local matrix multiplications. While SUMMA-like algorithms have the advantages of being straightforward to implement and generally scaling well, for sparse problems, they have the disadvantage of forcing processes to proceed in lockstep in the inner loop of the algorithm. When there is load imbalance, either in the amount of computation that must be performed locally by each process, the amount of data that must be transferred to each process, or both, this can potentially result in decreased performance. This can occur even when the total amount of work performed by each process is the same, since processes may have differing amounts of work in each iteration. To illustrate, Figure 1 shows two kinds of load balance for squaring a randomized sparse matrix generated by the R-MAT model (Chakrabarti et al., 2004), with parameters $a=0.6$, $b=c=d=0.4/3$, edgefactor $8$, and scale $17$. The total (end-to-end) computation has only $20\%$ load imbalance whereas the synchronization points in between stages amplify the load imbalance to 2.3$\approx$ 2.3$\times$ 2.3 ×. The load imbalance is defined as the max/avg ratio, the ratio of maximum number of flops performed by any processor to the average number of flops per processor.
@@
Load imbalance is often addressed by randomly permuting the rows and/or columns of the sparse matrices, but this has disadvantages, the foremost being that the required permutation of the inputs, followed by permutation of the outputs, can be expensive. Furthermore, permuting the matrix may possibly remove structural locality, resulting in decreased performance of local matrix multiply operations. For example, Slota et al. (Slota et al., 2017) showed that the loss of locality due to random permutations can hurt the performance of graph algorithms significantly where parallel PageRank performs 25×2-5$\times$2 - 5 × slower on a randomly permuted graph compared to other orderings. Even if we ignore the loss of locality or try to recover it by local reordering after a global random permutation, there are theoretical limits on how much load balance random permutations can achieve. For matrices with dense rows or columns, the load balance can deteriorate with $\sqrt{p}$ for large $p$, where $p$ is the number of processors in the worst case, per Theorem 5.4 of Azad et al. (Azad et al., 2020).
@@
Figure 1. Total (end-to-end) vs.  per-stage load balance multiplying a R-MAT model-generated sparse matrix with a sparse 2D algorithm. Simulated on a $16\times 16$ process grid.
@@
2. Background
Sparse times dense matrix multiply (SpMM) and sparse times sparse matrix multiply (SpGEMM) are two important sparse linear algebra primitives. SpMM is used in a variety of blocked iterative methods, graph algorithms, as well as graph neural networks (GNNs) (Huang et al., 2020; Tripathy et al., 2020; Hu et al., 2020). In these contexts, SpMM typically involves multiplication of a sparse matrix with a tall and skinny dense matrix, where the number of columns of the sparse matrix varies between 32 and 1024. SpGEMM is also used in graph algorithms, including betweenness centrality, triangle counting, cycle detection, and Markov clustering, as well as in applications such as in genomics (Buluç and Gilbert, 2011; Azad et al., 2015; Dongen, 2000; Bustamam et al., 2012; Davis, 2018).
2.1. Distributed Matrix Algorithms
Distributed matrix algorithms allow multiple processes to work together to execute a single matrix multiply operation. Let’s first consider a distributed matrix multiply computing the product $C=AB$. Here, each of the matrices $C$, $A$, and $B$ is split up into tiles which live on different processes. The matrix tiles will be sparse or dense, depending on what kind of matrix multiplication is being performed. Let’s assume that $C$ is distributed among a grid of $M\times N$ tiles, $A$ is distributed among a grid of $M\times K$ tiles, and $B$ is distributed among a grid of $K\times N$ tiles. The algorithms we consider in this paper rely on a tiling scheme that is regular on the indices. In other words, each tile has the same dimensions, modulo differences due to the tile size not dividing the matrix size evenly. A different, more irregular tiling mechanism might be more appropriate for matrices from certain applications such as those involving linear-scaling quantum-chemical calculations (Borštnik et al., 2014), where the near-sightedness creates natural block-sparse structure. However, the vast majority of application targets (GNNs, graph algorithms, eigenvector computations, and tensor decompositions) do not have this natural block-sparse structure, so the added complexity of irregular tiling is not justified for us.
@@ 2.1. Distributed Matrix Algorithms
2.2. Bulk Synchronous SUMMA
@@
As discussed in Section 1, variation in the number of nonzeros, as well as the sparsity patterns, of individual tiles of sparse input matrices can create load imbalance in computation due to the differing numbers of flops that must be performed in each local matrix multiply, as well as load imbalance in communication, due to the different amounts of data that must be transferred by each process. Bulk synchronous SUMMA-like algorithms can suffer due to these load imbalances.
2.3. RDMA and Asynchrony
Remote Direct Memory Access (RDMA) is a hardware capability offered by the Network Interface Cards (NICs) in most modern supercomputers and datacenters. Using RDMA, a process can expose a region of its memory, often referred to as “shared memory,” to be directly accessible by remote processes. Processes can then issue put, get, and atomic operations to remotely access and manipulate the shared memory regions of other processes. These RDMA operations are handled directly by the NIC on the remote node, hence no coordination is required in order to access the memory of a remote process. Some of the commonly used frameworks that support direct RDMA access include OpenSHMEM, GASNet-EX, and MPI’s one-sided communication API. For distributed programs that run on GPUs, systems with Infiniband can currently take advantage of GPUDirect RDMA, which allows direct transfers from GPU-to-GPU over the network. NVSHMEM, an extension of OpenSHMEM from NVIDIA, allows for direct transfers of data between GPUs, either within a node over PCIe or NVLink, or between nodes using GPUDirect RDMA over Infiniband. In this paper, we use NVSHMEM to directly copy between GPUs. A number of different libraries have been developed that use RDMA primitives to build data structures and algorithms, with some focusing particularly on irregular data structures and algorithms, where asynchronous RDMA primitives can be particularly beneficial, both for performance and software engineering reasons (Brock et al., 2019; Bachan et al., 2017; Fürlinger et al., 2016).
3. RDMA-Based Algorithms
3.1. Data Structures
@@ 3.1.1. Reading Tiles
3.1.2. Modifying Remote Tiles
@@
3.2. Algorithms
3.2.1. RDMA Stationary C Algorithm
The most straightforward RDMA-based algorithm is the stationary C algorithm, in which each processor iterates through its tiles of the C matrix, and, for each tile, iterates through the corresponding row block of A and column block of B, retrieving each of the remote tiles via get_tile and multiplying them together using a local matrix multiply operation. Since the method get_tile does not require any coordination with remote processors, each process can execute its work independently, and does not need to wait on or synchronize with other processors. In order to achieve good performance, a few optimizations are required, including prefetching tiles for the next iteration, which allows overlap of computation and communication, and an iteration offset in the inner loop, which balances which tile is requested within each row and column and generally ensures the first RDMA get is to the local tile. These optimizations are discussed in detail in Section 3.3.
{minted}
[fontsize=]pythonfor i in 0..M-1: for k in 0..K-1: if A.owner(i, k) == rank(): for j in 0..N-1: local_a = A.tile_ref(i, k) local_b = B.get_tile(k, j) local_c = local_a*local_b queue[C.owner(i, j)].push(get_ptr(local_c))
while local_c_ptr = queue[rank()].pop(): C.my_tile() += get_tile(local_c_ptr) barrier()
@@ 3.2.1. RDMA Stationary C Algorithm
3.2.2. RDMA Stationary A and B Algorithms
@@ 3.3. Optimizations
{minted}
[fontsize=]pythonfor i in 0..M-1: for j in 0..N-1: if C.owner(i, j) == rank(): k_offset = i + j buf_a = A.async_get_tile(i, k_offset buf_b = B.async_get_tile(k_offset for k_in 0..K-1: k = (k_+ k_offset) local_a = buf_a.get() local_b = buf_b.get() local_c = C.tile_ref(i, j)
if (k_+ 1 < K): buf_a = A.async_get_tile(i, (k+1) buf_b = B.async_get_tile((k+1) local_c += local_a*local_b barrier()
@@ 3.4. Workstealing Algorithms
In locality-aware workstealing, instead of stealing randomly, each process will only steal pieces of work for which it owns one of the components, ensuring a lower cost of performing stolen work. For example, in a stationary C locality-aware work-stealing algorithm, processes will first attempt to perform matrix multiplies involving their tiles of C. After these are completed, each process will attempt to steal work involving their tiles of A and B. Locality-aware workstealing requires a 3D work grid, where element $i,j,k$ corresponds to the component matrix multiply $C[i,j]+=A[i,k]*B[k,j]$. There is a slightly higher overhead associated with this 3D work grid, which requires a separate remote fetch-and-add for each piece of work performed, but this is largely offset by the improvement in locality.
{minted}
[fontsize=]pythondef attempt_work(i, k): # Remote atomic fetch-and-add to reserve work my_j = reserve_grid[i, k]++ while my_j < N: local_a = A.get_tile(i, k) local_b = B.get_tile(k, j) local_c = local_a*local_b
queue[C.owner(i, j)].push(get_ptr(local_c)) my_j = reserve_grid[i, k]++
# Do work for my tiles for i in 0..M-1: for k in 0..K-1: if A.owner(i, k) == rank(): attempt_work(i, k)
# Attempt to steal work for idx in 0..M-1: i = (rank()+idx) / M k = (rank()+idx) attempt_work(i, k)
while local_c_ptr = queue[rank()].pop(): C.my_tile() += get_tile(local_c_ptr) barrier()
@@ 4. Performance Model
If we offset the order of iteration of the $k$ loop by $i+j$, as discussed in Section 3.3, we can also ensure that communication will be balanced with each process sending exactly two tiles per iteration, one of the A matrix, and one of the B matrix. To prove this, assume tile $(i,j)$ is stored on processor $i(N+1)+j$, where $i\in[0,N)$, $j\in[0,N)$, and $N=\sqrt{p}$.
@@
Figure 2. Inter-node roofline plots for SpMM and SpGEMM with a 2D distribution. SpMM plot models performance for different widths of the dense B matrix at a fixed scale (24 GPUs), while SpGEMM models performance at different scales. Dashed horizontal lines represent local roofline peaks for SpMM and SpGEMM operations, while vertical lines represent inter-node roofline peaks for particular problems.
@@
Note that the roofline model is dependent on the sparsity structure, since $cf$ depends on the particular sparse matrices being multiplied. Our inter-node SpGEMM roofline model is plotted in Figure 2. To obtain realistic values for cf and the number of flops performed FLOPS(A, B), we performed each of the component local SpGEMM operations in the distributed SpGEMM operation for the isolates subgraph4 matrix, recording their values for cf and FLOPS(A, B) to compute average values for different numbers of GPUs $P$. In the SPGEMM plot, inter-node roofline peaks are much closer to their local roofline peaks than in the SpMM plot. Thus, our roofline model suggests SpGEMM is significantly less network communication-bound than SpMM, although still communication bound. We discuss more conclusions from our roofline model in Section 6.
5. Implementation
5.1. Communication Layer
As discussed in Section 2.3, the current generation of GPU-based supercomputers offers GPUDirect RDMA over Infiniband networks, which ensures that issuing one-sided operations will be efficiently executed in hardware. In GPUDirect RDMA, the CPU prepares an Infiniband request, which it sends to the local network interface card (NIC). When the NIC receives the request, it will copy the data directly over the network from GPU to GPU, without staging data through the CPU. When copies take place between GPUs within the same node, the high-bandwidth NVLink fabric is used.
@@ 5.2. Data Structures
We implement sparse and dense distributed matrix data structures using BCL, an RDMA-based distributed data structures library in written in C++. Both dense and sparse matrix data structures split the distributed matrix up into evenly sized tiles, which are then assigned to processors using a processor grid, as in ScaLAPACK. As discussed in Section 3.1, each process has a copy of a directory of remote pointers, which point to the matrix tiles stored in pinned memory, directly accessible through RDMA operations using NVSHMEM. In the dense case, a single remote pointer points to a dense tile data structure, and in the sparse case three remote pointers point to the values, row pointer, and column index arrays of a compressed sparse row (CSR) data structure. All data is stored directly on GPUs, and can be copied directly from GPU to GPU over the network using BCL’s NVSHMEM backend.
5.3. Data Access Primitives